OpenStudy (anonymous):

Which function has x-intercepts of 3 and 8 and contains the point (1, 2)? (Points : 1)

OpenStudy (anonymous):

I think it is A but I'm confused

OpenStudy (anonymous):

x intercept would be the point for which Y=0. all the functions given have x=3,and x=8 as x-intercept. put x=1 and y=2 in all of the equations for the equation whose right and left sides would be equal is the required result you are looking for. for example if i put x=1 and y=2 in first equation \[\Large 2=\frac{-1}{36}(1-3)(1-8)\] this gives \[\Large 2=\frac{14}{36}\] \[\Large 2=\frac{7}{18}\] since 2 is not equal to 7/18 so its not the answer. do the same for remaining options for which you get 2=2 would be the proper answer.

OpenStudy (anonymous):

@LeilaJudeh did you get it?

OpenStudy (anonymous):

Yea, thank you very much!